OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL The position of an object at time t is given by s(t) = -2 - 6t. Find the instantaneous velocity at t = 2 by finding the derivative.

Nnesha (nnesha):

replace t by 2 and then solve

OpenStudy (anonymous):

-14?

Nnesha (nnesha):

yes

OpenStudy (anonymous):

that's it?

Nnesha (nnesha):

hmm i guess find derivative

Nnesha (nnesha):

i guess nope*

OpenStudy (solomonzelman):

instantaneous velocity at x=a is f'(a).

OpenStudy (solomonzelman):

yes, you need the derivative. And then to plug in 2 for t.

OpenStudy (anonymous):

how do i get the derivative?

OpenStudy (solomonzelman):

differentiate (find derivative of) of each term. What is the derivative of -6x? What is the derivative of -2?

OpenStudy (anonymous):

idk how to find it

OpenStudy (solomonzelman):

do you know the power rule? \(\large\color{slate}{ \displaystyle \frac{d}{dx} (x^n)=n\cdot x^{n-1} }\)

OpenStudy (solomonzelman):

(d/dx is a notation for a derivative)

OpenStudy (solomonzelman):

Have seen this rule before?

OpenStudy (solomonzelman):

I will give you an example: \(\large\color{slate}{ \displaystyle \frac{d}{dx} (x^3)=3\cdot x^{3-1} =3x ^2 }\)

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \frac{d}{dx} (x^1)=? }\)

OpenStudy (anonymous):

1*x^0 =1

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

okay so how do i use that for the question?

OpenStudy (solomonzelman):

Now, there is a different rule, with a constant C. \(\large\color{slate}{ \displaystyle \frac{d}{dx} (c~x^n)=c \cdot \left( \frac{d}{dx}x^n \right)=c\cdot \left( n\cdot x^{n-1} \right) }\)

OpenStudy (solomonzelman):

you are basically taking the constant out.

OpenStudy (solomonzelman):

For example, \(\large\color{slate}{ \displaystyle \frac{d}{dx} (3~x^4)=3 \cdot \left( \frac{d}{dx}x^4 \right)=3\cdot \left(4\cdot x^{4-1} \right) = 3\cdot \left(4 x^{3} \right)=12x^3 }\)

OpenStudy (solomonzelman):

You need, \(\large\color{slate}{ \displaystyle \frac{d}{dx} (-6~x^1)= }\)

OpenStudy (solomonzelman):

(I took out a 3, and YOU are taking -6 out)

OpenStudy (anonymous):

i am still confused about how i take the -6 out

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \frac{d}{dx} (-6~x^1)=(-6) \cdot \left( \frac{d}{dx}x^1 \right)=~~....}\)

OpenStudy (solomonzelman):

see?

OpenStudy (solomonzelman):

lost?

OpenStudy (anonymous):

yep

OpenStudy (solomonzelman):

\(\color{slate}{ \displaystyle \frac{d}{dx} (c~x^n)=c \cdot \left( \frac{d}{dx}x^n \right)=c\cdot \left(n\cdot x^{n-1} \right) = c\cdot \left(n x^{n-1} \right)=cnx^{n-1} }\)

OpenStudy (solomonzelman):

Example, \(\large\color{slate}{ \displaystyle \frac{d}{dx} (-2~x^6)=(-2) \cdot \left( \frac{d}{dx}x^6 \right)=(-2)\cdot \left(6\cdot x^{6-1} \right) = (-2)\cdot \left(6 x^{5} \right) \\ ~\\~\\=-18x^{5} }\)

OpenStudy (solomonzelman):

do you see what I am doing in the example?

OpenStudy (solomonzelman):

(I am separating the -2, and the derivative of x^6)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{d}{dx}\left( -6x^{^{_{\Large 1}}} \right)=(-6)\frac{d}{dx}\left( x^{^{_{\Large 1}}} \right) }\)